# Agentic Search in the Wild: Intents and Trajectory Dynamics from 14M+ Real Search Requests

### v1

@@ Abstract.
LLM-powered search agents are increasingly being used for multi-step information seeking tasks, yet the IR community lacks empirical understanding of how agentic search sessions unfold and how retrieved evidence is used. This paper presents a large-scale log analysis of agentic search based on 14.44M search requests (3.97M sessions) collected from DeepResearchGym, i.e. an open-source search API accessed by external agentic clients. We sessionize the logs, assign session-level intents and step-wise query-reformulation labels using LLM-based annotation, and propose Context-driven Term Adoption Rate (CTAR) to quantify whether newly introduced query terms are traceable to previously retrieved evidence. Our analyses reveal distinctive behavioral patterns. First, over 90% of multi-turn sessions contain at most ten steps, and 89% of inter-step intervals fall under one minute. Second, behavior varies by intent. Fact-seeking sessions exhibit high repetition that increases over time, while sessions requiring reasoning sustain broader exploration. Third, agents reuse evidence across steps. On average, 54% of newly introduced query terms appear in the accumulated evidence context, with contributions from earlier steps beyond the most recent retrieval. The findings suggest that agentic search may benefit from repetition-aware early stopping, intent-adaptive retrieval budgets, and explicit cross-step context tracking. We plan to release the anonymized logs to support future research.
@@ 1. Introduction
We plan to release an anonymized version of the logs to support reproducibility (details in Section 3).
@@ 3.1. DRGym Log Overview
To support reproducibility and enable further research, we will release the cleaned and anonymized logs associated with this study. Prior to release, we will remove direct identifiers (e.g., IP addresses) and apply standard PII scrubbing on free-text fields, releasing only the fields needed to reproduce our analyses with anonymized session IDs. This release has already received formal approval from the DRGym organizers. We will document the anonymization procedure and residual risks in the dataset card, following prior large-scale LLM interaction log releases and privacy-preserving analyses (Handa et al., 2025; Zheng et al., 2024). After the initial publication, the dataset will be updated as additional logs are collected and validated.

### v2

@@ Abstract.
LLM-powered search agents are increasingly being used for multi-step information seeking tasks, yet the IR community lacks empirical understanding of how agentic search sessions unfold and how retrieved evidence is used. This paper presents a large-scale log analysis of agentic search based on 14.44M search requests (3.97M sessions) collected from DeepResearchGym, i.e. an open-source search API accessed by external agentic clients. We sessionize the logs, assign session-level intents and step-wise query-reformulation labels using LLM-based annotation, and propose Context-driven Term Adoption Rate (CTAR) to quantify whether newly introduced query terms are traceable to previously retrieved evidence. Our analyses reveal distinctive behavioral patterns. First, over 90% of multi-turn sessions contain at most ten steps, and 89% of inter-step intervals fall under one minute. Second, behavior varies by intent. Fact-seeking sessions exhibit high repetition that increases over time, while sessions requiring reasoning sustain broader exploration. Third, agents reuse evidence across steps. On average, 54% of newly introduced query terms appear in the accumulated evidence context, with contributions from earlier steps beyond the most recent retrieval. The findings suggest that agentic search may benefit from repetition-aware early stopping, intent-adaptive retrieval budgets, and explicit cross-step context tracking. We have released the anonymized logs to support future research and reproducibility, available at a public Hugging Face repository.
@@ 1. Introduction
We have released the anonymized logs to support future research and reproducibility, available at this public Hugging Face repository.
@@ 3.1. DRGym Log Overview
To support reproducibility and enable further research, we have released the cleaned and anonymized logs associated with this study at https://huggingface.co/datasets/cx-cmu/deepresearchgym-agentic-search-logs. We have removed direct identifiers (e.g., IP addresses) and applied standard PII scrubbing on free-text fields, releasing only the fields needed to reproduce our analyses with anonymized session IDs. This release has received formal approval from the DRGym organizers. We have documented the anonymization procedure and residual risks in the dataset card, following prior large-scale LLM interaction log releases and privacy-preserving analyses (Handa et al., 2025; Zheng et al., 2024). After the initial publication, the dataset may be updated as additional logs are collected and validated.